A train leaves the station at noon traveling at 44 mph. One hour later, another train going 52 mph, travels north on a parallel track. How many hours will the second travel before it overtakes the first train?

please help me...it will be appreciated.

Answer by ptaylor(2198)

Distance(d) equals Rate(r) times Time(t) or d=rt; r=d/t and t=d/r
Let t=number of hours that elapses before the second train catches the first (counting from when the second train leaves)
Distance travelled by the first train=44+44t (had 1 hour head start)
Distance travelled by the second train=52t
Now we know that when the above two distances are equal, the second train will have caught up with the first, so:
44+44t=52t subtract 44t from each side
44+44t-44t=52t-44t collect like terms
44=8t divide each side by 8
t=5.5 hours-------------------------Number of hours the second train travels before it overtakes the first
CK
44+44*5.5=52*5.5
44+242=286
286=286

Hope this helps---ptaylor